OpenStudy (ksaimouli):

integrate 3x^2e^2x dx

OpenStudy (ksaimouli):

\[\int\limits_{}^{} 3x^2e^2x dx\]

OpenStudy (ksaimouli):

i took \[u=X^2 dv=e^2x\]

OpenStudy (ksaimouli):

\[du=2x dx v=\frac{ e^2x }{ 2 }\]

OpenStudy (ksaimouli):

\[uv-\int\limits\limits_{}^{}v dv\]

OpenStudy (ksaimouli):

\[(\frac{ 3 }{ 2 }e ^{2x} x^2 -\]

OpenStudy (ksaimouli):

after that confused

OpenStudy (ksaimouli):

how to integrate v

OpenStudy (anonymous):

\[\int3x^2e^{2x}dx?\\ 3\int x^2e^{2x}dx\] First, integrate by parts, letting \[\begin{matrix}u=x^2& &dv=e^{2x}dx\\ du=2x\;dx& &v=\frac{1}{2}e^{2x}\end{matrix}\\ 3\int x^2e^{2x}dx=3\left[\frac{1}{2}x^2e^{2x} - \frac{1}{2}\int2xe^{2x}dx\right]\] Now, I'd let \[t=2x\\ dt=2\;dx\\ \frac{1}{2}dt=dx\] \[3\int x^2e^{2x}dx=\frac{3}{2}x^2e^{2x} - \frac{3}{2}\int te^tdt\] Integrate by parts again.

OpenStudy (ksaimouli):

thx

OpenStudy (ksaimouli):

can u show me how to integrate sin(x^3) i mean just tell me how to do i will to the rest

OpenStudy (ksaimouli):

@SithsAndGiggles

OpenStudy (anonymous):

There is no closed form of that integral. The best you can do is find an approximation to the function sin(x³), then integrate the approximation over a given interval.

OpenStudy (ksaimouli):

does this include series because i am seeing for the first time is this cal 2

OpenStudy (ksaimouli):

tylors series

OpenStudy (anonymous):

You could do that, yes. I think there's a way using the power series for that function, but I'm not so sure about that. Do you know how to find the Taylor polynomial of the function?

OpenStudy (ksaimouli):

noo i have no idea may be we will learn at the end of march next chapter but i am curious to know if u have free time u could explain or else ur wish

OpenStudy (anonymous):

I'll post a link to a video that explains it pretty well, if you're patient enough to watch it: https://www.khanacademy.org/math/calculus/sequences_series_approx_calc/maclaurin_taylor/v/maclauren-and-taylor-series-intuition I've only watched the older version, so I'm not sure how much better the updated one may be.

OpenStudy (ksaimouli):

ya thx i will check that out but ^ how did u get (3rd step) 2xe^2x dx

OpenStudy (anonymous):

That's \[v\;du = \frac{1}{2}2xe^{2x}dx,\text{ with the $\frac{1}{2}$ factored out.}\]

OpenStudy (ksaimouli):

i got that i mean it is \[\int\limits_{}^{}v \] so v=\[\frac{ 1 }{ 2 }e^2x\]

OpenStudy (ksaimouli):

i dont get where u got 2x in front of e^2x (typo ^ it is e^(2x))

OpenStudy (anonymous):

The 2x comes from the du term. Since u = x², we get du = 2x dx

OpenStudy (ksaimouli):

ohh so it is \[\int\limits_{}^{}v du\]

OpenStudy (anonymous):

Yep, integration by parts, using the u and dv substitutions, yields \[uv-\int v\;du\]

OpenStudy (ksaimouli):

okay

OpenStudy (ksaimouli):

thx

OpenStudy (anonymous):

You're welcome